OpenStudy (anonymous):

Trigonometric Proof Help? \[1 + \cot (x)=\cos(x)(\sec(x)+\csc(x))\]

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hello!

OpenStudy (misty1212):

most all of this is algebra, very little is trig

OpenStudy (anonymous):

Yeah, I honestly understand most of this, I just have trouble starting equations :/

OpenStudy (misty1212):

we can work with the right hand side and turn it in to the left hand side in one step

OpenStudy (misty1212):

what you need to know is that \(\sec(x)=\frac{1}{\cos(x)}\) and also that \(\csc(x)=\frac{1}{\sin(x)}\)

OpenStudy (anonymous):

So just change the identities and solve from there?

OpenStudy (misty1212):

yes, you get it right away \[\cos(x)(\sec(x)+\csc(x))=\cos(x)\left(\frac{1}{\cos(x)}+\frac{1}{\sin(x)}\right)\]

OpenStudy (anonymous):

So common denomonator between 1/sinx and 1/cosx

OpenStudy (misty1212):

no don't add, just multiply out on the right

OpenStudy (anonymous):

Oh! Thats actually very convenient

OpenStudy (misty1212):

isn't it just?

OpenStudy (misty1212):

you are done right?

OpenStudy (anonymous):

yup! thank you!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]